OpenStudy (4meisu):

x + 3y = 7 and -x + 4y = 7 Solve using linear combinations (elimination)

OpenStudy (anonymous):

x + 3y = 7 -x + 4y = 7 => 7y=14=>y=2 which on subs in 1st eq gives x=1... got it ??

OpenStudy (4meisu):

yes, i think i do

OpenStudy (anonymous):

okay